OpenStudy (anonymous):

PLEASE HELP!!! I will give metal and if u r correct, i will FAN! I don't have much time left.

OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to \[Y=\frac{ 1 }{ 4 }X+7\] and that pass through the point (-8, 9).

OpenStudy (solomonzelman):

reverse the slope saying that if it is A, then the perp will be -1/a Can you write the slope for me? The write (with whatever A) y=Ax+ (unknown y-intercept "b") plug in (-8,9) -8 for x and 9 for y, and solve for y-intercept (knowing the slope from the first step).

OpenStudy (solomonzelman):

I am not very good at explaining -:(

OpenStudy (anonymous):

slope is 1/4

OpenStudy (solomonzelman):

yes, but in your case it would be \(\Large\color{blue}{ \bf -\frac{1}{\frac{1}{4} } }\)

OpenStudy (anonymous):

why?

OpenStudy (solomonzelman):

ohh the equations of parallel and perp. So for the parallel one just say y=1/4 x + b plug in (-8,9) and solve for b and for the perp one y= -4 x +b plug in yur point (the same way) and solve for b

OpenStudy (solomonzelman):

Perpendicular slope to slope m is \(\Large\color{blue}{ \bf -\frac{1}{m} }\)

OpenStudy (anonymous):

that is not how i learned it...

OpenStudy (solomonzelman):

okay, good. You can do it.... I'll help you more if you need to.

OpenStudy (anonymous):

is it parallel: \[y=\frac{ 1 }{ 4 }x+11\] perpendicular: y = -4x - 23

OpenStudy (solomonzelman):

yes correct !!

OpenStudy (solomonzelman):

Good work

OpenStudy (anonymous):

sweet! thx

OpenStudy (solomonzelman):

My pleasure !